# Joint parameter estimations for spin glasses

### v1

@@ 2.1. Auxiliary Lemmas
$C_{1}$ $:=2\beta+\|\partial_{y}f\|_{\infty},$
$C_{2}$ $:=2\beta+5\beta^{2}+2\beta\|\partial_{y}f\|_{\infty}+\|\partial_{y}^{2}f\|_{\infty}.$
As an immediate corollary of this lemma with $f(x,y)=y$ and $f(x,y)=x$, we have
@@
Proof of Lemma 2.1.
Let distinct $i,j\in[N]$ be fixed. Set
$m_{j}^{(i)}({\sigma}):=\sum_{k\neq i,j}J_{jk}\sigma_{k}$
and $L^{(i)}_{j}({\sigma})$ is same as $L_{j}({\sigma})$ with $m_{j}({\sigma})$ replaced by $m_{j}^{(i)}({\sigma})$. Since $L_{i}(\tau)$ is the expectation of $f(\tau_{i},\tau_{i}m_{i}({\tau}))$ conditionally on $(\tau_{l})_{l\neq i}$, we have
$\bigl{\langle}\bigl{(}f(\tau_{i},\tau_{i}m_{i}({\tau}))-L_{i}({\tau})\bigr{)}\bigl{(}f(\tau_{j},\tau_{j}m_{j}({\tau}))-L_{j}^{(i)}({\tau})\bigr{)}\bigr{\rangle}=0,$
which implies
$\bigl{\langle}\bigl{(}f(\tau_{i},\tau_{i}m_{i}({\tau}))-L_{i}({\tau})\bigr{)}\bigl{(}f(\tau_{j},\tau_{j}m_{j}({\tau}))-L_{j}({\tau})\bigr{)}\bigr{\rangle}$
$=\bigl{\langle}c_{ij}({\tau})(f(\tau_{i},\tau_{i}m_{i}({\tau}))-L_{j}({\tau}))\bigr{\rangle},$
where
$c_{ij}({\tau}):=L_{j}^{(i)}({\tau})-L_{j}({\tau}).$
Now, using Taylor expansion,
$\bigl{|}L_{j}^{(i)}({\tau})-L_{j}({\tau})+J_{ij}\partial_{J_{ij}}L_{j}({\tau})\bigr{|}$ $\leqslant\frac{J_{ij}^{2}}{2}\max_{J_{ij}\in\mathbb{R}}\bigl{|}\partial_{J_{ij}}^{2}L_{j}({\tau})\bigr{|}.$
Here,
$\partial_{J_{ij}}L_{j}({\tau})$ $=\frac{e^{\beta m_{j}({\tau})+h}\partial_{y}f(1,m_{j}({\tau}))+e^{-\beta m_{j}({\sigma})-h}\partial_{y}f(-1,-m_{j}({\tau}))}{e^{\beta m_{j}({\tau})+h}+e^{-\beta m_{j}({\tau})-h}}$
$+\beta\frac{e^{\beta m_{j}({\tau})+h}f(1,m_{j}({\tau}))-e^{-\beta m_{j}({\tau})-h}f(-1,-m_{j}({\tau}))}{e^{\beta m_{j}({\tau})+h}+e^{-\beta m_{j}({\tau})-h}}$
$-\beta\frac{e^{\beta m_{j}({\tau})+h}f(1,m_{j}({\tau}))+e^{-\beta m_{j}({\tau})-h}f(-1,-m_{j}({\tau}))}{e^{\beta m_{j}({\tau})+h}+e^{-\beta m_{j}({\tau})-h}}\tanh(\beta m_{j}({\tau})+h)$
@@
$|\partial_{J_{ij}}L_{j}({\tau})|$ $\leqslant C_{1}:=2\beta+\|\partial_{y}f\|_{\infty}.$
Since $\partial_{J_{ij}}L_{j}({\tau})$ is independent of $i,$ it follows that
$\Bigl{|}\sum_{i\neq j}J_{ij}\partial_{J_{ij}}L_{j}({\tau})\Bigr{|}=\Bigl{|}\sum_{i,j=1}^{N}J_{ij}\partial_{J_{ij}}L_{j}({\tau})\Bigr{|}$ $\leqslant C_{1}N\|J\|.$
On the other hand, since
$|\partial_{J_{ij}}^{2}L_{j}({\tau})|$ $\leqslant C_{2}:=2\beta+5\beta^{2}+2\beta\|\partial_{y}f\|_{\infty}+\|\partial_{y}^{2}f\|_{\infty},$
we have
$\Bigl{|}\sum_{i\neq j}J_{ij}^{2}\partial_{J_{ij}}^{2}L_{j}({\tau})\Bigr{|}=\Bigl{|}\sum_{i,j=1}^{N}J_{ij}^{2}\partial_{J_{ij}}^{2}L_{j}({\tau})\Bigr{|}$
$\leqslant C_{2}\sum_{i,j=1}^{N}J_{ij}^{2}=C_{2}{\bf 1}^{T}(J\circ J){\bf 1}\leqslant C_{2}N\|J\circ J\|\leqslant C_{2}N\|J\|^{2}.$
From these, we arrive at
$\sum_{i\neq j}|c_{ij}({\tau})|\leqslant C_{1}N\|J\|+\frac{C_{2}}{2}N\|J\|^{2}$
and thus,
$\sum_{i\neq j}\bigl{\langle}\bigl{(}f(\tau_{i},\tau_{i}m_{i}({\tau}))-L_{i}({\tau})\bigr{)}\bigl{(}f(\tau_{j},\tau_{j}m_{j}({\tau}))-L_{j}({\tau})\bigr{)}\bigr{\rangle}$ $\leqslant C_{1}N\|J\|+\frac{C_{2}}{2}N\|J\|^{2}.$
This readily yields our assertion. $\blacksquare$
Let $0<\delta<M$ be fixed. Let $\rho$ be a nonnegative twice differentiable function on $\mathbb{R}$ with $\rho\equiv 1$ on $(\delta,\infty)$, $\rho\equiv 0$ on $(-\infty,0)$, and $\|\rho^{\prime}\|_{\infty},\|\rho^{\prime\prime}\|_{\infty}<\infty$. Also, let $\bar{\rho}$ be a nonnegative twice differentiable function on $\mathbb{R}$ with $\bar{\rho}\equiv 1$ on $(-\infty,-\delta)$, $\bar{\rho}\equiv 0$ on $(0,\infty)$, and $\|\bar{\rho}^{\prime}\|_{\infty},\|\bar{\rho}^{\prime\prime}\|_{\infty}<\infty.$ Define $f(x,y)=x\rho(y)$ and $\bar{f}(x,y)=x\bar{\rho}(y).$ Set
$\Delta(x,z)$ $=f(x,xz)-\frac{e^{\beta z+h}f(1,z)+e^{-\beta z-h}f(-1,-z)}{e^{\beta z+h}+e^{-\beta z-h}},$
$\bar{\Delta}(x,z)$ $=\bar{f}(x,xz)-\frac{e^{\beta z+h}\bar{f}(1,z)+e^{-\beta z-h}\bar{f}(-1,-z)}{e^{\beta z+h}+e^{-\beta z-h}}.$
Our next lemma establishes some useful bounds for the difference on the left-hand side of (14) associated to $f$ and $\bar{f}$ defined above.
Lemma 2.3.
We have that
$\Delta(1,z)\geqslant 0,\,\,\forall z\in\mathbb{R}\,\,\mbox{and}\,\,\Delta(1,z)\geqslant\frac{1}{1+e^{2\beta M+2h}},\,\,\forall\delta\leqslant z\leqslant M,$ (15)
$\Delta(-1,z)\geqslant 0,\,\,\forall z\in\mathbb{R}\,\,\mbox{and}\,\,\Delta(-1,z)\leqslant-\frac{1}{1+e^{-2\beta\delta-2h}},\,\,\forall\delta\leqslant z\leqslant M,$ (16)
$\bar{\Delta}(1,z)\geqslant 0,\,\,\forall z\in\mathbb{R}\,\,\mbox{and}\,\,\bar{\Delta}(1,z)\geqslant\frac{1}{1+e^{-2\beta\delta+2h}},\,\,\forall-M\leqslant z\leqslant-\delta,$ (17)
$\bar{\Delta}(-1,z)\leqslant 0,\,\,\forall z\in\mathbb{R}\,\,\mbox{and}\,\,\bar{\Delta}(-1,z)\leqslant-\frac{1}{1+e^{2\beta M-2h}},\,\,\forall-M\leqslant z\leqslant-\delta.$ (18)
Proof.
Computing directly yields that for any $z\in\mathbb{R},$
$\Delta(1,z)=\frac{\rho(z)+\rho(-z)}{1+e^{2\beta z+2h}}\geqslant 0,\quad\Delta(-1,z)=-\frac{\bar{\rho}(z)+\bar{\rho}(-z)}{1+e^{-2\beta z-2h}}\leqslant 0,$
$\bar{\Delta}(1,z)=\frac{\bar{\rho}(z)+\bar{\rho}(-z)}{1+e^{2\beta z+2h}}\geqslant 0,\quad\bar{\Delta}(-1,z)=-\frac{\bar{\rho}(z)+\bar{\rho}(-z)}{1+e^{-2\beta z-2h}}\leqslant 0.$
If $\delta\leqslant z\leqslant M,$ then
$\Delta(1,z)$ $=\frac{1}{1+e^{2\beta z+2h}}\geqslant\frac{1}{1+e^{2\beta M+2h}},$
$\Delta(-1,z)$ $=-\frac{1}{1+e^{-2\beta z-2h}}\leqslant-\frac{1}{1+e^{-2\beta\delta-2h}}.$
Also, if $-M\leqslant z\leqslant-\delta,$
$\bar{\Delta}(1,z)$ $=\frac{1}{1+e^{2\beta z+2h}}\geqslant\frac{1}{1+e^{-2\beta\delta+2h}},$
$\bar{\Delta}(-1,z)$ $=-\frac{1}{1+e^{-2\beta z-2h}}\leqslant-\frac{1}{1+e^{2\beta M-2h}}.$

### v2

@@ 2.1. Auxiliary Lemmas
$c_{ij}({\tau})$ $:=L_{j}^{(i)}({\tau})-L_{j}({\tau}),$
$X_{i}(\tau)$ $:=f_{i}(\tau_{i},m_{i}({\tau}))-L_{i}({\tau}).$
We bound (15) first. Note that
$|X_{i}(\tau)|\leqslant 2C(1+|m_{i}(\tau)|).$ (16)
For $t\in[0,1],$ let $J^{(i,j)}(t)$ be a symmetric square matrix with $J_{ab}^{(i,j)}(t)=J_{ab}$ for all $(a,b)\neq(i,j)$ and $(j,i)$ and $J_{ij}^{(i,j)}(t)=J_{ji}^{(i,j)}(t)=(1-t)J_{ij}.$ Denote $\mathcal{L}_{ij}(t)$ by the function $L_{j}(\tau)$ with $J$ being replaced by $J^{(i,j)}(t)$. Using Taylor’s formula yields that
$L_{j}^{(i)}({\tau})$ $=\mathcal{L}_{ij}(1)$
$=\mathcal{L}_{ij}(0)+\mathcal{L}_{ij}^{\prime}(0)+\frac{1}{2}\int_{0}^{1}(1-s)\mathcal{L}_{ij}^{\prime\prime}(s)ds$
$=L_{j}(\tau)+J_{ij}Y_{j}+J_{ij}^{2}Z_{j}^{(i)},$
where
$Y_{j}$ $=-\partial_{J_{ij}}L_{j}(\tau),$
$Z_{j}^{(i)}$ $=\frac{1}{2}\int_{0}^{1}(1-s)\partial_{J_{ij}}^{2}L_{j}(\tau)\Big{|}_{J=J^{(i,j)}(s)}ds.$
We emphasize that here $Y_{j}$ depends only on $j$ since
$\partial_{J_{ij}}L_{j}({\tau})$ $=\frac{e^{\beta m_{j}({\tau})+h}\partial_{y}f_{j}(1,m_{j}({\tau}))+e^{-\beta m_{j}({\sigma})-h}\partial_{y}f_{j}(-1,m_{j}({\tau}))}{e^{\beta m_{j}({\tau})+h}+e^{-\beta m_{j}({\tau})-h}}$
$+\beta\frac{e^{\beta m_{j}({\tau})+h}f_{j}(1,m_{j}({\tau}))-e^{-\beta m_{j}({\tau})-h}f_{j}(-1,m_{j}({\tau}))}{e^{\beta m_{j}({\tau})+h}+e^{-\beta m_{j}({\tau})-h}}$
$-\beta\frac{e^{\beta m_{j}({\tau})+h}f_{j}(1,m_{j}({\tau}))+e^{-\beta m_{j}({\tau})-h}f_{j}(-1,m_{j}({\tau}))}{e^{\beta m_{j}({\tau})+h}+e^{-\beta m_{j}({\tau})-h}}\tanh(\beta m_{j}({\tau})+h).$
Now from the given assumption,
$|\partial_{J_{ij}}L_{j}({\tau})|$ $\leqslant 2\beta C(1+|m_{j}(\tau)|)+\|\partial_{y}f_{j}\|_{\infty}=:K_{j}.$
Similarly one can compute the second derivative of $L_{j}({\tau})$ with respect to $J_{ij}$; the exact form is not important here, but it is crucial to realize that
$|\partial_{J_{ij}}^{2}L_{j}({\tau})|$ $\leqslant\tilde{C}\bigl{(}1+|m_{j}(\tau)|+\|\partial_{y}f_{j}\|_{\infty}+\|\partial_{y}^{2}f_{j}\|_{\infty}\bigr{)}$
for some constant $\tilde{C}$ depending only on $C$ and $\beta.$ It follows from these bounds that
$|Y_{j}|\leqslant K_{j}$ (17)
and
$|Z_{j}^{(i)}|$ $\leqslant\tilde{C}\Bigl{(}1+\int_{0}^{1}(1-s)|(J^{(i,j)}(s)\tau)_{j}\bigr{|}ds+\|\partial_{y}f_{j}\|_{\infty}+\|\partial_{y}^{2}f_{j}\|_{\infty}\Bigr{)}$
$\leqslant\tilde{C}\bigl{(}1+2\|J\|+\|\partial_{y}f_{j}\|_{\infty}+\|\partial_{y}^{2}f_{j}\|_{\infty}\bigr{)}=:\tilde{K}_{j},$ (18)
where we have used the bound $\|J^{(i,j)}(s)\|\leqslant\|J\|+(1-s)|J_{ij}|\leqslant 2\|J\|.$ Now, write
$\Big{|}\sum_{i\neq j}X_{i}(\tau)c_{ij}(\tau)\Big{|}\leqslant\Big{|}\sum_{i\neq j}J_{ij}X_{i}(\tau)Y_{j}(\tau)\Big{|}+\sum_{i\neq j}\frac{J_{ij}^{2}}{2}|X_{i}(\tau)||Z_{j}^{(i)}|.$
Note that $J_{ij}=0$ when $i=j$. Using (16) and (17) leads to
$\Bigl{|}\sum_{i\neq j}J_{ij}X_{i}(\tau)Y_{j}(\tau)\Bigr{|}$ $\leqslant 2C\|J\|\Bigl{(}\sum_{i=1}^{N}(1+|m_{i}(\tau)|)^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{j=1}^{N}K_{j}^{2}\Bigr{)}^{1/2}.$
Similarly, from (16) and (18) and using $\|J\circ J\|\leqslant\|J\|^{2},$
$\sum_{i\neq j}J_{ij}^{2}|X_{i}(\tau)||Z_{j}^{(i)}|$ $\leqslant C\|J\|^{2}\Bigl{(}\sum_{i=1}^{N}(1+|m_{i}(\tau)|)^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{i=1}^{N}\tilde{K}_{i}^{2}\Bigr{)}^{1/2}.$
From these, (15) is bounded by
$\Big{|}\sum_{i\neq j}X_{i}(\tau)c_{ij}(\tau)\Big{|}$ $\leqslant 2C\|J\|\Bigl{(}\sum_{i=1}^{N}(1+|m_{i}(\tau)|)^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{j=1}^{N}K_{j}^{2}\Bigr{)}^{1/2}$
$+C\|J\|^{2}\Bigl{(}\sum_{i=1}^{N}(1+|m_{i}(\tau)|)^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{i=1}^{N}\tilde{K}_{i}^{2}\Bigr{)}^{1/2}.$
As for (14), using the Taylor expansion yields
$\Bigl{|}f_{i}(\tau_{j},m_{j}({\tau}))-f_{i}(\tau_{j},m_{j}^{(i)}({\tau}))+J_{ij}\tau_{i}\partial_{y}f_{i}(\tau_{j},m_{j}({\tau}))\Bigr{|}$ $\leqslant\frac{1}{2}J_{ij}^{2}\|\partial^{2}_{y}f_{i}\|_{\infty}.$
Here, we can control similar to the above,
$\Bigl{|}\sum_{i\neq j}X_{i}(\tau)\bigl{(}f_{j}(\tau_{j},m_{j}({\tau}))-f_{j}(\tau_{j},m_{j}^{(i)})\bigr{)}\Bigr{|}$
$\leqslant\Big{|}\sum_{i,j}J_{ij}(X_{i}(\tau)\tau_{i})\big{(}\partial_{y}f_{j}(\tau_{j},m_{j}({\tau}))\big{)}\Big{|}+\frac{1}{2}\sum_{i,j}J_{ij}^{2}|X_{i}(\tau)|\|\partial^{2}_{y}f_{j}\|_{\infty}$
$\leqslant 2C\|J\|\Bigl{(}\sum_{i=1}^{N}(1+|m_{i}(\tau)|)^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{i=1}^{N}\|\partial_{y}f_{i}\|_{\infty}^{2}\Bigr{)}^{1/2}$
$+C\|J\|^{2}\Bigl{(}\sum_{i=1}^{N}(1+|m_{i}(\tau)|)^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{i=1}^{N}\|\partial_{y}^{2}f_{i}\|_{\infty}^{2}\Bigr{)}^{1/2}.$
Putting everything together and noting that $\sum_{i=1}^{N}m_{i}(\tau)^{2}\leqslant N\|J\|^{2}$, we have
$\sum_{i\neq j}\bigl{\langle}\bigl{(}f_{i}(\tau_{i},m_{i}({\tau}))-L_{i}({\tau})\bigr{)}\bigl{(}f_{i}(\tau_{j},m_{j}({\tau}))-L_{j}({\tau})\bigr{)}\bigr{\rangle}$
$\leqslant K(\|J\|+\|J\|^{2})\Bigl{(}N+N\|J\|^{2}+\sum_{i=1}^{N}\|\partial_{y}f_{i}\|_{\infty}^{2}+\sum_{i=1}^{N}\|\partial_{y}^{2}f_{i}\|_{\infty}^{2}\Bigr{)}$
for some constant $K$ depending only on $\beta$ and $C.$ Finally, for the sum over $i=j$, it can be easily bounded by $C^{\prime\prime}N(+\|J\|^{2})$ for some constant $K^{\prime}$ depending only on $C$ and this completes our proof.
$\blacksquare$
As an immediate corollary of this lemma with $f(x,y)=xy$ and $f(x,y)=x$, we have
@@
Now we start with the proof of Theorem 1.1 existence part.
2.2. Proof of Theorem 1.1:Existence of the MPLE
We begin with the following lemma, which identifies sufficient conditions for the existence of MPLE.
Lemma 2.3.
Let $\sigma\in\{-1,+1\}^{N}$ and assume that there exist distinct $1\leqslant i,j,k,\ell\leqslant N$ and $a\in\mathbb{R}$ such that
$\begin{split}&\sigma_{i}=1,m_{i}({\sigma})>a,\,\,\sigma_{j}=-1,m_{j}({\sigma})>a,\\ &\sigma_{k}=1,m_{k}({\sigma})<a,\,\,\sigma_{l}=-1,m_{l}({\sigma})<a.\end{split}$ (19)
then $\lim_{|\beta|+|h|\to\infty}\bar{L}(\beta,h|\sigma)=-\infty.$
Proof.
For brevity of notation, we denote $m_{s}(\sigma)$ by $m_{s}$. Noting that
$\bar{L}(\beta,h|{\sigma})$ $=\frac{1}{N}\sum_{s=1}^{N}\Big{(}\sigma_{s}(\beta m_{s}+h)-\log\cosh(\beta m_{s}+h)-\log 2\Big{)}$
$\leqslant\frac{1}{N}\sum_{s=1}^{N}\Big{(}|\beta m_{s}+h|-\log\cosh(\beta m_{s}+h)-\log 2\Big{)}$
and each term in the above sum is non-positive, it is enough to show that given $\sigma$}, for any $K>0$, there exists some $C>0$ such that the following statement holds: for any $(\beta,h)\in\mathbb{R}^{2}$ with $|\beta|+|h|\geqslant C,$ there exists some $s=s(\beta,h)\in[N]$ such that
$\sigma_{s}(\beta m_{s}+h)\leqslant-K.$ (20)
We will argue by contradiction. Assume that $i,j,k,$ and $l$ satisfy (19) and each of them violates (20). Then we must have
$\sigma_{i}=1,\quad\beta m_{i}+h>-K$
$\sigma_{j}=-1,\quad\beta m_{j}+h<K$
$\sigma_{k}=1,\quad\beta m_{k}+h>-K$
$\sigma_{l}=-1,\quad\beta m_{l}+h<K,$
This confines $h$ to the interval:
$\max(-\beta m_{i}-K,-\beta m_{k}-K)<h<\min(-\beta m_{j}+K,-\beta m_{l}+K).$ (21)
For this interval to be non-empty, its upper bound must be greater than its lower bound.
Case $\beta\geqslant 0$: The conditions $m_{j}>a>m_{k}$ imply $m_{j}>m_{k}$. The non-empty interval constraint implies $\beta m_{k}-K<-\beta m_{j}+K$, which simplifies to $\beta(m_{j}-m_{k})<2K$. Therefore, we have
$\beta<2K/(m_{j}-m_{k}).$ (22)
Case $\beta<0$: The conditions $m_{i}>a>m_{l}$ imply $m_{i}>m_{l}$. The constraint implies $\beta m_{i}-K<-\beta m_{l}+K$, which simplifies to $\beta(m_{i}-m_{l})<2K$. Therefore, we have
$\beta<2K/(m_{i}-m_{l}).$ (23)
Combining (22) and (23), we obtain
$|\beta|<\max\left(\frac{2K}{m_{j}-m_{k}},\frac{2K}{m_{i}-m_{l}}\right),$
which, in combination with (21) forces $|h|$ to be bounded. So, $|\beta|+|h|$ can not be arbitrarily large. This concludes the proof of the lemma. $\blacksquare$
The following lemma, which is purely deterministic, states a key step to verify the conditions in Lemma 2.3. Let
$\tilde{T}_{N}:=N^{-2}\sum_{i,j=1}^{N}\operatorname{sech}^{2}(\beta_{0}m_{i}+h_{0})\operatorname{sech}^{2}(\beta_{0}m_{j}+h_{0})(m_{i}-m_{j})^{2}.$
Note that this is a deterministic object and different from (8).
Lemma 2.4.
Let $\beta_{0}>0,h_{0}\in\mathbb{R},$ and $c>0$. There exist positive constants $\delta,\epsilon,K$, depending only on $\beta_{0},h_{0}$ and $c$, such that for any sequence $m_{1},m_{2},\ldots,m_{N}$ of real numbers satisfying $\tilde{T}_{N}\geqslant c$, there exists an integer $r$ with $|r|\leqslant K/\delta$ and $K\leqslant(r-1)\delta<r\delta\leqslant K$ such that
$|\{i\in[N]:m_{i}\in[-K,(r-1)\delta]\}|\geqslant\varepsilon N\text{ and }|\{i\in[N]:m_{i}\in[r\delta,K]\}|\geqslant\varepsilon N.$ (24)
Proof.
Since $\operatorname{sech}^{2}(\beta_{0}x+h_{0})\leqslant\min(1,4e^{-2(\beta_{0}|x|-|h_{0}|)})$, we can choose $K>0$ sufficiently large so that
$\sum_{i,j:|m_{i}|,|m_{j}|\leqslant K}(m_{i}-m_{j})^{2}\geqslant\frac{c}{2}N^{2},$ (25)
which implies that at least $\eta N$ of the $m_{i}$’s lie within [K,K][-K,K], where $\eta=\sqrt{c/2}/(2K)$. Set
$\varepsilon=\min\left(\frac{\eta}{4},\frac{c}{80K^{2}}\right)\,\,\text{and}\,\,\delta=\min\left(\frac{\sqrt{c}}{5},\frac{K}{2}\right)$
and define
$r=\max\left\{s\in\mathbb{Z}:\left|\left\{i\in[N]:m_{i}\in[s\delta,K]\right\}\right|\geqslant\varepsilon N\right\}.$
Suppose, for contradiction, that
$\left|\left\{i\in[N]:m_{i}\in[-K,(r-1)\delta]\right\}\right|<\varepsilon N.$
This implies that at most $2\varepsilon N$ of the $m_{i}$ can lie in the interval $K,K]\setminus[(r-1)\delta,(r+1)\delta]$. We now derive an upper bound on the left-hand side of (25).
$(I)$ The total contribution from pairs $i,j$ such that at least one of $m_{i}$ or $m_{j}$ lies outside $(r-1)\delta,(r+1)\delta]$ is at most
$2\cdot N\cdot 2\varepsilon N\cdot 4K^{2}\leqslant\frac{c}{5}N^{2}.$
$(II)$ The total contribution from pairs $i,j$ such that both $m_{i}$ and $m_{j}$ lie within $(r-1)\delta,(r+1)\delta]$ is at most
$N^{2}\cdot(2\delta)^{2}\leqslant\frac{c}{5}N^{2}.$
By the choice of $\varepsilon$ and $\delta$, the sum of these contributions is at most $2cN^{2}/5$, which contradicts (25). Therefore,
$\left|\left\{i\in[N]:m_{i}\leqslant(r-1)\delta\right\}\right|\geqslant\varepsilon N$
and the lemma follows. $\blacksquare$
Now we complete the proof of the existence part of Theorem 1.1.
Proof.
For $\ell,u\in\mathbb{R}$ with $\ell<u$ let
$S_{\ell,u}:=\left\{i\in[N]:m_{i}\in[\ell,u]\right\}.$
First of all, we claim that for any $a\in\mathbb{R}$ and $\epsilon,\delta>0,$ the following events have negligible probability as $N\to\infty,$
$|S_{-K,a}|\geqslant\varepsilon N\,\,\text{and}\,\,\mbox{the spins in}\,\,\{\sigma_{i}:i\in S_{-\infty,a+\delta/3}\}\text{ have all the same signs}$ (26)
@@
$|S_{a+\delta,K}|\geqslant\varepsilon N\,\,\mbox{and}\,\,\mbox{the spins in}\,\,\{\sigma_{i}:i\in S_{a+2\delta/3,\infty}\}\text{ have all the same signs}.$ (27)
We will only prove (27) since (26) can be treated similarly. Suppose $|S_{a+\delta,K}|\geqslant\varepsilon N$ and $\sigma_{i}=1$ for all $i\in S_{a+2\delta/3,\infty}$. Choose a smooth function $\rho:\mathbb{R}\to[0,1]$ such that $\rho\equiv 1$ on $a+\delta,\infty)$ and $\rho$ vanishes outside $a+2\delta/3,\infty).$ Take $f(\sigma_{i},m_{i})=\sigma_{i}\rho(m_{i}).$ Note that for any $i\in[N],$
$f(\sigma_{i},m_{i})-\langle f(\sigma,m_{i})\rangle_{-i}$ $=\rho(m_{i})\bigl{(}\sigma_{i}-\tanh(\beta m_{i}+h)\bigr{)}.$
This is nonnegative if $i\in S_{a+2\delta/3,\infty}$ and is equal to zero if $i\notin S_{a+2\delta/2,\infty}.$ On the other hand, there exists a constant $\kappa>0$ such that
$f(\sigma_{i},m_{i})-\langle f(\sigma_{i},m_{i})\rangle_{-i}\geqslant\kappa\text{ for }i\in S_{a+\delta,K}.$
Together with the assumption $|S_{a+\delta,K}|\geqslant\varepsilon N$, we have
$\sum_{i=1}^{N}\big{(}f(\sigma_{i},m_{i})-\langle f(\sigma_{i},m_{i})\rangle_{-i}\big{)}\geqslant\kappa\varepsilon N,$
contradicting the concentration results established in Lemma 2.1. Hence, (27) must be negligible and this completes the proof of our claim. Next, let $\eta\in(0,1)$ be fixed. From the positivity assumption of $\widetilde{T}_{N}$ and Lemma 2.4, there exist positive constants, $\delta,\epsilon,K$ depending on $\beta_{0},h_{0},$ and $c$ such that for any $N$ large enough, with probability at least $1-\eta$, there exists some $a\in\mathbb{R},$
$|S_{-K,a}|\geqslant\epsilon N\,\,\mbox{and}\,\,|S_{a+\delta,K}|\geqslant\epsilon N$
Noting that there are only finitely many possible $a$’s; indeed, $a=(r-1)\delta$ for some integer $r\in\mathbb{Z}$ with $|r|\leqslant K/\delta$. It follows from the above claim and an union bound argument that as long as $N$ is large enough, with probability at least $1-2\eta,$ the condition (19) in Lemma 2.3 hold and thus, on this event, $\lim_{|\beta|+|h|\to\infty}\bar{L}(\beta,h|\sigma)=-\infty.$